What is 24 C in room atmospheric pressure?
Answer #1

Atmospheric pressure = P = 101325 * (°C + 273) / 273 Pascals

P = 101325 * (24 + 273)/273 = 110232.7 Pascal OR 110232.7/101325 = 1.088 atm .....Answer